The induced magnetic field at radial distance 7.4 mm from the central axis of a circular parallel-plate capacitor 2.2 × 10-6 T. The plates have radius 4.0 mm. At what rate is the magnitude of the electric field between the plates changing?
Answer #1

criven B = 2-2x10-61 radius rs 7-4X10-3m ro = 4X10-3 m B = usko vo* x de 28 2.2x10-6 - 42x107x8.85 x 1042 x(24X10-3) 2 X 764X